What is the apparent mass of a 1-kg object that has been accelerated to 90.8% of light speed?

Answer in units of kg.
Answer #1

Apparent mass is given by:

m = m0*gamma

gamma = lorentz factor = 1/sqrt (1 - v^2/c^2)

v = 0.908*c

gamma = 1/sqrt (1 - 0.908^2) = 2.3868

So,

m = m0*gamma = 1*2.3868

m = apparent mass = 2.3868 kg